A student is taking a​ multiple-choice exam in which each question has two choices. Assuming that she has no knowledge of the correct answers to any of the​questions,she has decided on a strategy in which she will place two balls​( marked (A and B) into a box. She randomly selects one ball for each question and replaces the ball in the box. The marking on the ball will determine her answer to the question. There are six ​multiple-choice questions on the exam. Complete parts​(a)through​(d)below.

b.What is the probability that she will get at least five questions​ correct?

_____

​(Roundto four decimal places as​needed.)
Answer #1

P(getting correct answer), p = 0.5

Number of questions, n = 6

q = 1 - p = 0.5

Binomial distribution: P(X) = nCx px qn-x

P(she will get at least five questions​ correct) = P(5) + P(6)

= 6C5x0.55x0.5 + 0.56

= 0.1094